Steel Young's Modulus (Y) 20 X 1010 N/m2 Shear Modulus (S) 8.1 X 1010 N/m2 Bulk...

Steel
Young's Modulus (Y) 20 X 1010 N/m2
Shear Modulus (S) 8.1 X 1010 N/m2
Bulk Modulus (B) 16 X 1010 N/m2

The table to the above represents various properties of steel. You have steel wire 4.4 meters in length that stretches 0.12 cm when subjected to a force of 300 N.

1)What is the diameter of the wire?

d1 =__________ cm

3)What would the diameter of the wire be if you wanted the wire to stretch 0.06 cm less when subjected to this same force?

d2 =______________ cm
Answer #1

1) let the diameter of the wire be 'd' m

now, stress = force/cross-sectional area = 300/(pi*d2/4) = 1200/(pi*d2)

now, as per Hooke's Law,

stress = young's moldulus*strain

strain = change in length/original length = 0.12/440 = 2.73*10-4

thus,  1200/(pi*d2) = (20 X 1010)*(2.73*10-4) = 2.45*107

or, d2 = 7*10-6 m

or, d = 2.646*10-3 m

or, diameter = 0.2646 cm

2) strain = 0.06/440 = 1.364*10-4

thus, 1200/(pi*d2) = (20 X 1010)*(1.364*10-4) = 2.73*107

or, d2 = 1.4*10-5 m

or, d = 3.74*10-3 m

or diameter = 0.374 cm